OpenStudy (anonymous):

Solving Systems of Equations by Substitution. -5x + y = -2 -3x + 6y =-12

OpenStudy (anonymous):

Well y=5x-2 So -3x+5x-2=-12, solve

OpenStudy (anonymous):

Multiply the top by -6 and add the equations together, then solve for x. Plug the x back in to get y

OpenStudy (anonymous):

paintballcs's method is actually more pure substitution, so I would do it that way

OpenStudy (anonymous):

y = -2 + 5x -3x + 6 (-2 + 5x) = -12 -3x -12 + 30x = -12 27x = 0 x = 0